TIME-RESOLVED INFRARED DIODE LASER SPECTROSCOPY OF THE ν_{1} BAND OF THE FeNO RAD-ICAL PRODUCED BY THE ULTRAVIOLET LASER PHOTOLYSIS OF Fe(CO)_{2}(NO)_{2}

\underline{SEIKI IKEDA}, MOTOKI NAKASHIMA and KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Rovibrational transitions of the ν_{1} band (N-O stretch) of the FeNO radical were observed in the 1750-1780 cm^{−1} region. The FeNO radical was produced by 193 nm excimer laser photolysis of Fe(CO)_{2}(NO)_{2} and the transient absorption signal was detected by time-resolved infrared diode laser spectroscopy. More than 40 lines were assigned to the ν_{1} fundamental band of the Ω = 5/2 spin component, together with ten Q-branch lines (J = 2.5 - 11.5) in the 1760 cm^{−1} region, to confirm the electronic ground state to be X^{2}Δ_{i}. Effective molecular constants for the Ω = 5/2 spin component, including the band origin ν_{0} (1767.26093(38) cm^{−1}), the rotational constant B (4610.17754(93) MHz) and the centrifugal distortion constant D (1.17003(47) kHz), were derived from a least squares fitting of the observed transitions. The average bond length r_{Co−N} between Co and N was calculated to be 1.621 Å from the rotational constant B_{0} assuming r_{N−O} =1.186 Å as given by ab initio calculation^{a}. The ν_{1} hot band lines originated from the ν_{2} (Fe-N-O bending; 308 cm^{−1} ^{a}) vibrationally excited state were also observed. Pure rotational lines of FeNO were also observed by the millimeter wave spectroscopy with the UV photolysis of Fe(CO)_{2}(NO)_{2}^{b}.

aM. Zhou and L. Andrews, J.P hys.Chem.A, 104, 3915 (2000)

bMicrowave session in this symposium.